A washer and a dryer cost $802 combined. The washer costs $48 less than the dryer. What is the cost of the dryer?
Answer by ikleyn(52794)

.

Let the dryer costs x dollars (an unknown value).


Then the washer costs (x-48) dollars, according to the condition.


Now I can write this condition

    dryer + washer = 802

as an equation

    (x-48) + x = 802.


Simplify and solve

    x - 48 + x = 802

    2x = 802 + 48

    2x = 850

     x = 850/2 = 425  dollars.


ANSWER.  The dryer costs 425 dollars.

Solved and explained in all details.

Is everything clear to you ?